A 1500-kg submarine rises straight up towards the surface. exerts both an upward buoyant force of 16160 the submarine, as well as a downward resistive force of 1130 N. What is the submarines aoceleration? Number Units
Answer #1

forces acting on the submarine are:

upward force =16160 N

downward resistive force=1130 N

downward weight=1500*9.8=14700 N

net force is along upward direction

net force magnitude=16160-1130-14700=330 N

then acceleration=net force/mass

=330/1500=0.22 m/s^2